OpenStudy (anonymous):

Help with algebra problem! Will medal and fan. 12x−2y−3+3(x+y)

OpenStudy (anonymous):

\[12x^{-2}y ^{-3}+3(x+y) when x = 2 and y = -3 so, 12(2)^{-2}(-3)^{-3}+3(2+-3)\]

OpenStudy (phi):

so the problem is to "evaluate" \[ 12(2)^{-2}(-3)^{-3}+3(2+-3) \] First, you need to know "order of operations" second, you should know this rule for negative exponents: \[ a^{-b}= \frac{1}{a^b} \] (flip and change the sign of the exponent ) can you simplify this part: 3(2+ -3) ? what is that ?

OpenStudy (anonymous):

Go to Mathway.com. They are really helpful.

OpenStudy (anonymous):

@phi, would the final answer be 6-9... -3?

OpenStudy (anonymous):

or would it be \[-3\div12(4)(9) = -3\div432\]

OpenStudy (phi):

just the 3(2-3) part is 6 -9 = -3 (or you could do 2-3 first = -1 and 3*-1= -3 ) that means you have , so far, \[ 12(2)^{-2}(-3)^{-3} - 3\] now lets do just \[ 2^{-2} \] can you re-write that using the flip and minus the exponent rule?

OpenStudy (anonymous):

\[12(-3)^-3-34\[12(-3^{3}) -3\div4\]

OpenStudy (anonymous):

Is this right? 12(-3)^-3 -3 /4

OpenStudy (phi):

\[ 2^{-2}= \frac{1}{2^2}= \frac{1}{2\cdot 2}= \frac{1}{4} \]

OpenStudy (phi):

ok with just that step?

OpenStudy (anonymous):

Yes. Makes sense.

OpenStudy (phi):

so far, we have \[ 12(2)^{-2}(-3)^{-3}+3(2+-3) \\ 12(2)^{-2}(-3)^{-3} - 3 \\ 12\cdot\frac{1}{4} \cdot (-3)^{-3} - 3 \] we replaced the 2^-2 with 1 /4

OpenStudy (phi):

now can you do \[ (-3)^{-3} \] using the flip and negate the exponent rule?

OpenStudy (anonymous):

Quick question - on my worksheet the last part is written (x+y)3 or (2+-3)3 once the variables are changed - does that still result in the same answer?

OpenStudy (phi):

is that (x+y)^3 or (x+y)*3 ?

OpenStudy (anonymous):

So that would be 1/9

OpenStudy (anonymous):

On my worksheet it is just a regular three. So unless it is a typo it is just (x+y)3

OpenStudy (phi):

if it's multiply for (x+y)*3 then we are ok. But usually if we multiply we put the numbers in front of the letters... but for the moment let's ignore that. back to (-3)^ -3 that is \[ \frac{1}{(-3)^3} = \frac{1}{-3 \cdot -3 \cdot -3} \]

OpenStudy (anonymous):

1/-27?

OpenStudy (anonymous):

so 12(1/4)(1/-27)-3

OpenStudy (phi):

yes. so far we have \[ 12\cdot\frac{1}{4} \cdot\frac{1}{-27} - 3 \\ \] notice we can simplify 12/4 to 3 to get \[ 3 \cdot\frac{1}{-27} - 3 \]

OpenStudy (phi):

or \[ -\frac{3}{27} - 3 \]

OpenStudy (phi):

what is -3/27 ? it is -1/9

OpenStudy (anonymous):

Yes so would the final answer be -1/9 - 3

OpenStudy (phi):

which is , as a mixed number \[ \frac{-1}{9}-\frac{27}{9} = - \frac{28}{9} \]

OpenStudy (phi):

you might also write it as \[ - 3 \frac{1}{9} \] negative 3 and one-nineths

OpenStudy (phi):

or -3.111....

OpenStudy (anonymous):

Thank you :)

OpenStudy (phi):

ok, I hope the steps made sense.

OpenStudy (anonymous):

They did!